OpenStudy (winsteria):

Please help if you can or just explain it really fast and easy What is the value of d? 11d = –44 d =???

OpenStudy (anonymous):

11d = -44 /11 /11 d = -4

OpenStudy (anonymous):

solve for d 11d=-44 divide both sides by 11 d=-44/11 d=-4

OpenStudy (winsteria):

That's what I thought

OpenStudy (winsteria):

Thank you both SO much

OpenStudy (anonymous):

I turned @sephiroth_07 to yellow!

OpenStudy (anonymous):

@SexyBeast890 yay? I am new here, what does that mean?

OpenStudy (anonymous):

Just your SS or smart score is higher :)

OpenStudy (anonymous):

cool tks